# A Phase-field Model for Apoptotic Cell Death

### v2

@@ 1. Introduction
Apoptosis, the process of programmed cell death, is a key component in a range of processes including cellular turnover, robust immune response to pathogens, embryonic development and chemical-induced cell death [elmore2007apoptosis]. Irregularities in this process, resulting in either enhanced or reduced apoptosis, are a causative factor in many human diseases, such as autoimmune and neurodegenerative disorders (excessive cell loss precipitating conditions such as Alzheimer’s disease) and many types of cancer (inadequate cell clearance) [bertheloot2021necroptosis, castillo2021apoptosis, vitale2023apoptotic]. Hence, the ability to modulate and understand the life or death of a cell is of high interest. Apoptotic cells can be identified by their distinct morphological features. These include separation from other cells, chromatin condensation and DNA fragmentation, irregular bulging in the plasma membrane (finger formation), cell shrinkage, and the formation of apoptotic bodies, which are then degraded in lysosomes [nossing202350, alberts2002programmed]. These morphological changes can be observed using microscopy techniques, with electron microscopy providing superior cellular details and reproducibility compared to light microscopy [kari2022programmed]. The highly complex and sophisticated mechanisms of apoptosis can be triggered by a range of factors, including components of the immune system and cytotoxic molecules administered as part of a chemotherapeutic regimen [elmore2007apoptosis]. Given the complexity of apoptosis, a thorough understanding of how environmental cues shape this process could provide a valuable platform for devising targeted therapies and refining diagnostic strategies.
Apoptosis can be initiated by one of two canonical pathways, see for instance, Lossi [Lossi2022_apoptosis].
@@
In this work, we refer to these mechanisms as intrinsic and extrinsic apoptosis, respectively. Irrespective of the initiating route, apoptosis involves a spatially distributed biochemical activation field that drives the progressive breakdown of cytoskeletal and organelle structures. Mathematically, this activation field is modelled in an agnostic manner, without distinguishing whether its origin is intrinsic or extrinsic. This activation field regulating death is referred to as the apoptotic phase, while the field representing the cell itself is denoted the cyto phase.
Moreover, according to Mustafa et al. [Mustafa2024_ApoptosisOverview], the extrinsic pathway begins when death ligands, such as TNF-$\alpha$ or Fas ligand, bind to their specific receptors on the cell membrane. In contrast, the intrinsic pathway involves permeabilisation of the outer mitochondrial membrane and the subsequent release of apoptogenic factors, including cytochrome c, under the control of Bcl-2 family proteins. Both routes lead to a cascade of biochemical reactions that activate caspases, fragment DNA, and promote the systematic disassembly of cellular structures.
@@
In this study, we develop a multi-component phase-field model incorporating biochemical reactions between a cyto phase (cell) and an apoptotic phase (triggered by the intrinsic or extrinsic apoptosis mechanism), to understand and model the physical characteristics exhibited by a cell undergoing apoptosis. This employs a coupled set of partial differential equations for the phase fields $\varphi$ (cyto phase) and $\sigma$ (apoptotic phase). Phase-field modelling of apoptosis remains a relatively unexplored area of modelling for cellular dynamics and therefore provides an opportunity for this model to serve as a novel tool for investigating apoptosis. The use of multi-component phase-field models with reaction equations is well-established, with extensive literature not only in general contexts [bai2022chemo, gaziano2024phase, moure2018three], but also in applications to biological phenomena, with published work to date focusing predominantly on modelling cell growth [wise2008three, hawkins2012numerical, colli2017optimal]. To obtain a chemo-mechanical phase-field model to produce the desired dynamics of finger formation, we take inspiration from Kobayashi’s work on solidification [kobayashi1993modeling]. Furthermore, to ensure the reaction component of the apoptosis model stays within certain bounds, we utilise the same substitution as Kobayashi [kobayashi1993modeling]. The apoptosis model is also inspired by the system proposed by Fried & Gurtin [fried1993continuum].
We address the need for a multi-component phase-field model for simulating apoptotic behaviour by developing a novel framework. The framework builds upon the mathematical structure for modelling reaction kinetics developed by Krambeck et al. [krambeck1970mathematical], which examines the fundamental properties of constitutive laws in chemically reacting systems, focusing on common physical assumptions and mass-action-based rate expressions. It also calls on the mechanics of phase-field models proposed by Clavijo et al. [clavijo2022continuum]. Additionally, further work by Clavijo et al. [clavijo2021extended] is also used, as it presents a unified, thermodynamically consistent framework for modelling multicomponent, solid systems undergoing coupled diffusion and chemical reactions. In their work, Clavijo et al. developed the extended Larché–Cahn framework to simulate how chemo-mechanical and reactive processes drive phase separation, domain evolution, and microstructural changes.
Numerical simulations illustrating finger formation (membrane blebs), nucleation and fragmentation are implemented in the Dedalus package [burns2020dedalus]. This package allows for rapid mathematical development through the symbolic implementation of the system into the numerical solver. Furthermore, it allows for temporal and spatial analysis in 1, 2, and 3-dimensional domains with minimal alterations to pre-existing code for higher-dimensional simulations.
@@
To demonstrate the model’s capabilities, we provide a comparison of numerical simulations at selected time points and their similarity with real-world experimental data (electron microscopy images from You et al. [you2015kml001]) in two morphological categories: membrane blebbing (or finger formation) and size and number of cavity formations (throughout cavity formations will be referred to as nucleations). In this manuscript, we describe nucleations as a void ($\varphi$=0) surrounded by the phase field $\varphi=1$ and fragments as an area occupied by $\varphi=1$ and surrounded by a void ($\varphi$=0). While variability of biological systems does result in significant differences between images, we demonstrate that there are consistent similarities between the model output and cell imaging.
@@
In Section §2, we develop a chemo-mechanical phase-field model to describe the dynamics of the cyto phase $\varphi$ and the apoptotic phase fields $\sigma$, including the system that produces features observed in biological cells undergoing apoptosis. In Section §3, we present computational simulations, used to illustrate finger formation, nucleation, and cell fragmentation during the degradation of the cyto phase field $\varphi$. Furthermore, in Section §4, we compare finger formation, nucleation, and cell fragmentation with electron microscopy images from You et al. [you2015kml001]. As mentioned, continuum models for apoptosis remain largely unexplored, presenting a unique opportunity not only to extend the utility of these models to a new biological context, but also to gain mechanistic insight into the interplay between chemical signalling and mechanical deformation during programmed cell death. In Section In Appendix §A, we present the mathematical derivation of the underlying configurational mechanics, including configurational forces and the configurational balance law.
@@ 2. A continuum theory for apoptotic processes
Although the precise molecular genesis of apoptosis remains an open question in biology, our model purposefully abstracts this uncertainty by focusing on the downstream enzymatic disassembly phase. The coordinated degradation of cytoskeletal and organelle structures while remaining agnostic to whether the apoptotic activation originates from intrinsic or extrinsic pathways.
To this end, consider the existence of two kinematic processes given by the phase fields $\varphi$ and $\sigma$. The process $\varphi$ shrinks and fades due to the presence of the process $\sigma$; therefore, we refer to $\varphi$ as the cyto phase field and $\sigma$ as the apoptotic phase field, triggering the enzymatic disassembly. To mimic the interplay between the cyto and apoptotic phases resulting in apoptosis, we adopted the framework from Clavijo et al. [Cla21] to model reaction-driven annihilation, where the phase fields $\varphi$ and $\sigma$ are treated as the concentrations of two species [^1] Although we treat the phase fields as concentrations when dealing with chemical reactions, these are mere representations of the underlying cyto and apoptotic agents., denoted by $\mathcal{A}_{\varphi}$ and $\mathcal{A}_{\sigma}$, with $\varphi\coloneq[\mathcal{A}_{\varphi}]$ and $\sigma\coloneq[\mathcal{A}_{\sigma}]$, the reaction describing these interactions is
(1) $\varsigma^{r}_{\varphi}\mathcal{A}_{\varphi}+\varsigma^{r}_{\sigma}\mathcal{A}_{\sigma}\stackrel{{\scriptstyle k}}{{\rightharpoonup}}\varsigma^{p}_{\varphi}\mathcal{A}_{\varphi},$
where $\varsigma^{r}_{\varphi}$, $\varsigma^{r}_{\sigma}$, and $\varsigma^{p}_{\varphi}$ are the stoichiometric coefficients for the reactants and products, respectively, and $k$ is the reaction rate constant.
2.1. Variational chemo-mechanical formulation
Here, we introduce a degenerate irreversible reaction model that describes a chemical-initiated enzymatic disassembly of the cyto phase field initiated by the apoptotic phase field. This degenerate irreversible reaction is given by
(2) $r\coloneqq\,\hat{k}g(\varphi)\,\sigma^{\varsigma^{r}_{\sigma}},\quad\text{with}\quad g(\varphi)=\varphi(1-\varphi),$
where $g(\varphi)$ promotes the development of the reaction at the interface; hence the degeneracy. Thus, the resulting internal mass supply rates for phase fields $\varphi$ and $\sigma$ are then
@@ 2.1. Variational chemo-mechanical formulation
where $(\varsigma_{\varphi}^{p}-\varsigma_{\varphi}^{r})r$ represents the net production or depletion rate of the (cyto) phase field $\varphi$, while $\varsigma_{\sigma}^{r}r$ corresponds to the vanishing rate of the apoptotic phase field $\sigma$ at the expense of the enzymatic disassembly of the cyto phase field $\varphi$. Depending on the values of the stoichiometric coefficients, the evolution of the phase field $\varphi$ can lead to either an increase or a decrease in the (cyto) phase field $\varphi$. For example, if $\varsigma_{\varphi}^{p}>\varsigma_{\varphi}^{r}$, then the phase field $\varphi$ increases, signifying growth, while $\varsigma_{\varphi}^{p}<\varsigma_{\varphi}^{r}$ leads to a net depletion of the phase field $\varphi$. We propose the following coupled gradient flows to model the kinematic interplay between these processes, mimicking chemically induced apoptotic dynamics.
Next, we consider the following energy functional
@@
where the free-energy density $\psi$ is given as a function of the two kinematic processes and their derivatives, namely
@@
where $f$ is a potential function depending on the phase field $\varphi$, and $\boldsymbol{\Sigma}$ is a symmetric positive definite matrix. The functional above splits the free energy into a potential function $f$, which may not be convex, and a regularisation term depending upon the gradients of the phase fields $\varphi$ and $\sigma$.
@@
where $\dot{\varphi}=\frac{\partial\varphi}{\partial t}$ and $\dot{\sigma}=\frac{\partial\sigma}{\partial t}$. Expression (7) describes the evolution of the phase fields $\varphi$ and $\sigma$. The parameters that dictate the relaxation towards equilibrium are $\tau>0$ for the phase field $\varphi$, and $\alpha>0$ and $\beta>0$ for $\sigma$. The first variation of $\delta\Psi$ with respect to the phase fields $\varphi$ and $\sigma$ are, respectively, given by
@@ 2.2. Thermodynamic potential
The polynomial structure is chosen to create a double-well potential. This setup drives phase separation, shaping an energy landscape that enables transitions between metastable states.
@@ 2.3. Specialisation to directional apoptosis
To formulate a directional form of the system (9) for simulations, we consider the anisotropic evolution in the phase field $\varphi$. We introduce the anisotropy into Equations (9) through the free energy $\psi$ in the term for $\boldsymbol{\Sigma}$, therefore we make $\boldsymbol{\Sigma}$ the following
(11) $\boldsymbol{\Sigma}=\begin{bmatrix}\dfrac{\epsilon^{2}}{2}&0\\ 0&\dfrac{\gamma^{2}}{2}\end{bmatrix}$
Let $\epsilon$ be a function of the gradient of the phase field $\varphi$. Thus, yielding the following system to describe directional anisotropy
(12) $\left\{\begin{aligned} \tau\dot{\varphi}&=-\left(\frac{\partial f}{\partial\varphi}-\mathrm{div}\mskip 2.0mu\left(\epsilon^{2}\nabla\varphi-\epsilon\frac{\partial\epsilon}{\partial\theta}\nabla\times(\varphi\boldsymbol{e}_{z})\right)\right)+(\varsigma_{\varphi}^{p}-\varsigma_{\varphi}^{r})r,\\[4.0pt] \alpha\dot{\sigma}&=\mathrm{div}\mskip 2.0mu\left(\gamma^{2}\nabla\sigma\right)+\beta\dot{\varphi}-\varsigma_{\sigma}^{r}r,\\[4.0pt] \epsilon&=\tilde{\epsilon}(\vartheta+\cos(\theta)),\\[4.0pt] \dfrac{\partial\epsilon}{\partial\theta}&=-\tilde{\epsilon}\vartheta\sin(\theta),\\[4.0pt] \theta&=\arctan\left(\frac{\partial_{y}\varphi}{\partial_{x}\varphi}\right),\\[4.0pt] r&=g(\varphi)\hat{k}\sigma^{\varsigma^{r}_{\sigma}},\end{aligned}\right$.
where $\partial_{v_{i}}\varphi=\frac{\partial\varphi}{\partial v_{i}}$ denotes partial derivatives along coordinates $v_{i}$. This system models directional growth dynamics through an anisotropic diffusion term influenced by the gradient of the phase field $\varphi$, with $\epsilon(\theta)$ introducing spatial dependencies. To maintain numerical stability during simulations, we make an approximation for $r=g(\varphi)\hat{k}\sigma^{\varsigma^{r}_{\sigma}}$ of the following form
(13) $r=g(\varphi)\dfrac{k_{1}}{\pi}\arctan(k_{2}\sigma)$

### v3

@@ 1. Introduction
Apoptosis, the process of programmed cell death, is a key component in a range of processes including cellular turnover, robust immune response to pathogens, embryonic development and chemical-induced cell death [8]. Irregularities in this process, resulting in either enhanced or reduced apoptosis, are a causative factor in many human diseases, such as autoimmune and neurodegenerative disorders (excessive cell loss precipitating conditions such as Alzheimer’s disease) and many types of cancer (inadequate cell clearance) [3, 5, 23]. Hence, the ability to modulate and understand the life or death of a cell is of high interest. Apoptotic cells can be identified by their distinct morphological features. These include separation from other cells, chromatin condensation and DNA fragmentation, irregular bulging in the plasma membrane (finger formation), cell shrinkage, and the formation of apoptotic bodies, which are then degraded in lysosomes [21, 1]. These morphological changes can be observed using microscopy techniques, with electron microscopy providing superior cellular details and reproducibility compared to light microscopy [15]. The highly complex and sophisticated mechanisms of apoptosis can be triggered by a range of factors, including components of the immune system and cytotoxic molecules administered as part of a chemotherapeutic regimen [8]. Given the complexity of apoptosis, a thorough understanding of how environmental cues shape this process could provide a valuable platform for devising targeted therapies and refining diagnostic strategies for personalised medication.
Apoptosis can be initiated by one of two canonical pathways, see for instance, Lossi [18].
@@
In this work, we refer to these mechanisms as intrinsic and extrinsic apoptosis, respectively. Irrespective of the initiating route, apoptosis involves a spatially distributed biochemical activation field that drives the progressive breakdown of cytoskeletal and organelle structures. Mathematically, this activation field is modelled in an agnostic manner, without distinguishing whether its origin is intrinsic or extrinsic. This activation field regulating death is referred to as the cytotoxic phase, while the field representing the cell itself is denoted the cyto phase.
Moreover, according to Mustafa et al. [20], the extrinsic pathway begins when death ligands, such as TNF-$\alpha$ or Fas ligand, bind to their specific receptors on the cell membrane. In contrast, the intrinsic pathway involves permeabilisation of the outer mitochondrial membrane and the subsequent release of apoptogenic factors, including cytochrome c, under the control of Bcl-2 family proteins. Both routes lead to a cascade of biochemical reactions that activate caspases, fragment DNA, and promote the systematic disassembly of cellular structures.
@@
In this study, we develop a multi-component non-conserved phase-field framework incorporating cascading biochemical reactions between a cyto phase (cell) and an apoptotic activation field (triggered by the intrinsic or extrinsic apoptosis mechanism), to understand and model the physical characteristics exhibited by a cell undergoing apoptosis. This employs a coupled set of partial differential equations for the fields $\varphi$ (cyto phase) and $\sigma$ (cytotoxic phase). In our scenario, $\sigma$ is a non-conserved external driving mechanism that triggers the degenerative reaction, which leads to apoptosis in the cell. The phase field $\varphi$ is also a non-conserved order parameter representing the cell. Phase-field modelling of apoptosis remains a relatively unexplored area of modelling for cellular dynamics and therefore provides an opportunity for this model to serve as a novel tool for investigating apoptosis. The use of multi-component phase-field models with reaction equations is well-established, with extensive literature not only in general contexts [2, 12, 19], but also in applications to biological phenomena, with published work to date focusing predominantly on modelling cell growth [24, 14, 7]. To obtain a phase-field model to produce the desired dynamics of finger formation, we take inspiration from Kobayashi’s work on solidification [16] and the system proposed by Fried & Gurtin [10].
@@
We address the need for a multi-component non-conserved phase-field framework for simulating apoptotic behaviour by developing a novel framework. To explore possible dynamics of this model, Numerical simulations illustrating finger formation (membrane blebs), nucleation and fragmentation are implemented in the Dedalus package [4]. To demonstrate the model’s capabilities, we provide a comparison of numerical simulations at selected time points and their similarity with real-world experimental data (electron microscopy images from You et al. [25]) in two morphological categories: membrane blebbing (or finger formation) and size and number of cavity formations (throughout cavity formations will be referred to as nucleations). In this manuscript, we describe nucleations as a void ($\varphi$=0) surrounded by the phase field $\varphi=1$ and fragments as an area occupied by $\varphi=1$ and surrounded by a void ($\varphi$=0). While variability of biological systems does result in significant differences between images, we demonstrate that there are consistent similarities between the model output and cell imaging. Furthermore, we present a possible methodology for a quantitative comparison, through the calculation of the area between the simulated cell $\varphi$ and the cell from the electron microscopy data. However, given the sparsity in the data we have at our disposal, a comparison is not made in this manuscript.
@@
In Section §2, we develop a phase-field framework to describe the dynamics of the cyto phase $\varphi$ and the cytotoxic phase fields $\sigma$, including the system that produces features observed in biological cells undergoing apoptosis. In Section §3, we present computational simulations, used to illustrate finger formation, nucleation, and cell fragmentation during the degradation of the cyto phase field $\varphi$. Furthermore, in Section §4, we compare finger formation, nucleation, and cell fragmentation with electron microscopy images from You et al. [25]. As mentioned, continuum models for apoptosis remain largely unexplored, presenting a unique opportunity not only to extend the utility of these models to a new biological context, but also to gain mechanistic insight into the interplay between chemical signalling and mechanical deformation during programmed cell death. In Appendix §A, we present the mathematical derivation of the underlying configurational mechanics, including configurational forces and the configurational balance law.
@@ 2. A continuum theory for apoptotic processes
Although the precise molecular genesis of apoptosis remains an open question in biology, our model purposefully abstracts this uncertainty by focusing on the downstream enzymatic disassembly phase. The coordinated degradation of cytoskeletal and organelle structures while remaining agnostic to whether the apoptotic activation originates from intrinsic or extrinsic pathways. In this work we consider a single cell interacting with its environment, on a fixed spatial domain $\Omega\subset\mathbb{R}^{2}$. Simulations focusing on many cells can be implemented through adjustment to the initial conditions, however, we only focus on single cell dynamics. Periodic boundary conditions are imposed on $\partial\Omega$ as we are not considering a closed system. To this end, consider the existence of two kinematic processes given by the phase fields $\varphi$ and $\sigma$. The process $\varphi$ shrinks and degrades due to the presence of the process $\sigma$.
To mimic the interplay between the cyto and cytotoxic phases resulting in apoptosis, we adopted the thermochemistry framework from Clavijo et al. [6] and the notation developed by Krambeck et al. [17] to model reaction-driven annihilation, where the phase fields $\varphi$ and $\sigma$ are treated as the concentrations of two species [^1] Although we treat the phase fields as concentrations when dealing with chemical reactions, these are mere representations of the underlying cyto and apoptotic agents., denoted by $\mathcal{A}_{\varphi}$ and $\mathcal{A}_{\sigma}$, with $\varphi\coloneq[\mathcal{A}_{\varphi}]$ and $\sigma\coloneq[\mathcal{A}_{\sigma}]$. The reaction describing these interactions are
(1) $\varsigma^{r}_{\varphi}\mathcal{A}_{\varphi}+\varsigma^{r}_{\sigma}\mathcal{A}_{\sigma}\stackrel{{\scriptstyle k_{1}}}{{\rightharpoonup}}\varsigma^{p}_{\varphi}\mathcal{A}_{\varphi},$
where $\varsigma^{r}_{\varphi}$, $\varsigma^{r}_{\sigma}$, and $\varsigma^{p}_{\varphi}$ are the stoichiometric coefficients for the reactants and products, respectively, and $k_{1}$ is the reaction rate constant and has the dimensions $1/T$. Given that this isn’t a specific chemical reaction the units for $k_{1}$ are determined through the duration of the experiment in which the comparison is made.
2.1. Phase-field framework formulation
Here, we introduce a degenerate irreversible reaction model that describes a chemical-initiated enzymatic disassembly of the cyto phase field initiated by the cytotoxic phase field. This degenerate irreversible reaction is given by
(2) $r\coloneqq\,k_{1}g(\varphi)\,\sigma^{\varsigma^{r}_{\sigma}},\quad\text{with}\quad g(\varphi)=\varphi(1-\varphi).$
The function $g(\varphi)$ promotes the development of the reaction at the interface between the cyto and cytotoxic phase. The resulting internal mass supply rates for phase fields $\varphi$ and $\sigma$ are
@@ 2.1. Phase-field framework formulation
where $(\varsigma_{\varphi}^{p}-\varsigma_{\varphi}^{r})r$ represents the net production or depletion rate of the cyto phase field $\varphi$, while $\varsigma_{\sigma}^{r}r$ corresponds to the vanishing rate of the cytotoxic phase field $\sigma$ at the expense of the enzymatic disassembly of the cyto phase field $\varphi$. Depending on the values of the stoichiometric coefficients, the evolution of the phase field $\varphi$ can lead to either an increase or a decrease in the cyto phase field $\varphi$. For example, if $\varsigma_{\varphi}^{p}>\varsigma_{\varphi}^{r}$, then the phase field $\varphi$ increases, signifying growth, while $\varsigma_{\varphi}^{p}<\varsigma_{\varphi}^{r}$ leads to a net depletion of the phase field $\varphi$. Thus, we propose the following coupled gradient flows to model the kinematic interplay between these processes, mimicking chemically induced apoptotic dynamics.
Next, consider the following energy functional
@@
where the free-energy density $\psi$ is given as a function of the cyto phase field and regularized by the gradients of both phase fields, namely
@@
where $f$ is a thermodynamical potential depending on the phase field $\varphi$ and $\boldsymbol{\Sigma}$ is a symmetric positive definite matrix. The functional above splits the free energy into a thermodynamical potential $f$, which may not be convex, and a regularisation term depending upon the gradients of the phase fields $\varphi$ and $\sigma$. The cytotoxic phase field $\sigma$ is not included in the free energy, as this is an external driving force for apoptosis and does not exhibit phase separation. Including $\nabla\sigma$ in the free energy acts as a regularisation term for the cytotoxic phase, this enforces diffusion as the main dynamic for the cytotoxic phase, as exhibited by real-world cytotoxic agents.
@@
where $\dot{\varphi}=\frac{\partial\varphi}{\partial t}$ and $\dot{\sigma}=\frac{\partial\sigma}{\partial t}$. The term $\beta\dot{\varphi}$ in the equation governing the evolution for $\sigma$ is included to create a latency state between the consumption of $\sigma$ and the structural degradation $\varphi$. The biological reasoning, for example, can be associated with the delay in the disassembly of the cytoskeleton or membrane permeabilisation. The Allen–Cahn type gradient flow is appropriately used in Equation (7), as apoptosis is an inherently non-conserved process, involving irreversible degradation and loss of cellular material. Expression (7) describes the evolution of the phase fields $\varphi$ and $\sigma$. The parameters that dictate the relaxation towards equilibrium are $\tau>0$ for the phase field $\varphi$, and $\alpha>0$ and $\beta>0$ for $\sigma$. The first variation of $\delta\Psi$ with respect to the phase fields $\varphi$ and $\sigma$ are, respectively, given by
@@ 2.2. Thermodynamic potential
The polynomial structure is chosen to create a double-well potential. This setup drives phase separation, shaping an energy landscape that enables transitions between stable states.
@@ 2.3. Specialisation to directional apoptosis
To formulate a directional form of the system (9) for simulations, we consider the anisotropic evolution in the phase field $\varphi$. The tensor $\boldsymbol{\Sigma}$ is assumed to be
(11) $\boldsymbol{\Sigma}=\begin{bmatrix}\dfrac{\epsilon^{2}(\theta)}{2}&0\\ 0&\dfrac{\gamma^{2}}{2}\end{bmatrix}$,
where $\epsilon$ is a function of the orientation $\theta$ of the interface, that is, the orientation of the gradient of the phase field $\varphi$. The dependency of $\epsilon$ on the orientation of the interface induces anisotropy during shrinkage, see for instance, the work by Kobayashi [16]. Thus, accounting for this dependency of $\epsilon$ yields the following system of equations
(12) $\left\{\begin{aligned} \tau\dot{\varphi}-(\varsigma_{\varphi}^{p}-\varsigma_{\varphi}^{r})r&=-\frac{\partial f}{\partial\varphi}+\mathrm{div}\mskip 2.0mu\left(\epsilon\frac{\partial\epsilon}{\partial\theta}\frac{\partial\theta}{\partial\nabla\varphi}|\nabla\varphi|^{2}+\epsilon^{2}\nabla\varphi\right),\\[4.0pt] \alpha\dot{\sigma}-\beta\dot{\varphi}+\varsigma_{\sigma}^{r}r&=\mathrm{div}\mskip 2.0mu(\gamma^{2}\nabla\sigma),\end{aligned}\right$.
where the variable $\theta$ is the orientation of the interface with respect to the horizontal axis and defined as
(13) $\theta=\arctan\left(\frac{\partial\varphi/\partial y}{\partial\varphi/\partial x}\right)$
Thus, the partial derivative of $\theta$ with respect to $\nabla\varphi$ yields
(14) $\frac{\partial\theta}{\partial\nabla\varphi}=\frac{1}{|\nabla\varphi|^{2}}\left(-\dfrac{\partial\varphi}{\partial y},\dfrac{\partial\varphi}{\partial x}\right)$
and noting that $\nabla\times\varphi\,\boldsymbol{e}_{z}=(-\partial\varphi/\partial y,\partial\varphi/\partial x)$, with $\boldsymbol{e}_{z}=[0,0,1]$, we arrive at
(15) $\left\{\begin{aligned} \tau\dot{\varphi}&=-\left(\frac{\partial f}{\partial\varphi}-\mathrm{div}\mskip 2.0mu\left(\epsilon^{2}\nabla\varphi-\epsilon\frac{\partial\epsilon}{\partial\theta}\nabla\times(\varphi\boldsymbol{e}_{z})\right)\right)+(\varsigma_{\varphi}^{p}-\varsigma_{\varphi}^{r})r,\\[4.0pt] \alpha\dot{\sigma}&=\mathrm{div}\mskip 2.0mu\left(\gamma^{2}\nabla\sigma\right)+\beta\dot{\varphi}-\varsigma_{\sigma}^{r}r,\\[4.0pt] \epsilon&=\tilde{\epsilon}(\vartheta+\cos(\theta)),\\[4.0pt] \dfrac{\partial\epsilon}{\partial\theta}&=-\tilde{\epsilon}\vartheta\sin(\theta),\\[4.0pt] \theta&=\arctan\left(\frac{\partial\varphi/\partial y}{\partial\varphi/\partial x}\right),\\[4.0pt] r&=k_{1}g(\varphi)\arctan(k_{2}\sigma).\end{aligned}\right$.
The parameter $\tilde{\epsilon}$ represents the interface thickness and $\vartheta$ is the degree of anisotropy, this parameter can be tuned depending on the number of protrusion exhibited during apoptosis. For numerical stability resasons, the reaction is approximated as shown in the system of Equations (15). The sensitivity of $\varphi$ to the cytotoxic phase field is controlled by the parameter $k_{2}$ and this is dimensionless, given that in this framework $\sigma$ is dimensionless. This system models directional growth dynamics through an anisotropic $\epsilon$ term influenced by the gradient of the phase field $\varphi$, with $\epsilon(\theta)$ introducing spatial dependencies. Shrinkage and growth in cells are anisotropic phenomena, leading to directional blebbing and fingering. In contrast, weaker apoptosis would result in more isotropic shrinkage.